OpenStudy (anonymous):

Graph the following linear system to determine the number of solutions. y equals negative three x plus six, six x plus two y equals twelve zero 1 2 infinitely many

OpenStudy (anonymous):

you just asked this didn't you ?

OpenStudy (anonymous):

yea but no has answered

OpenStudy (anonymous):

so if u know this can u help me plzzz

OpenStudy (anonymous):

I answered the last one....the answer is infinite solutions

OpenStudy (anonymous):

wat answer is that one for this one or the other one that is :Graph the following linear system to determine the number of solutions. y equals negative x plus two, x plus y equals five zero 1 2 infinitely many

OpenStudy (anonymous):

y = -2x + 2 --> rearrange to be 2x + y = 2 x + y = 5 now we have : 2x + y = 2 x + y = 5 -->(-1) --------------- 2x + y = 2 -x - y = -5 (result of multiplying by -1) ----------------add x = -3 now sub -3 in for x in either of the original equations x + y = 5 -3 + y = 5 y = 5 + 3 y = 8 check... 2x + y = 2 2(-3) + 8 = 2 - 6 + 8 = 2 2 = 2 (correct) This system of equations has 1 solution ... (-3,8)